# LangTraj: Diffusion Model and Dataset for Language-Conditioned Trajectory Simulation

### v1

@@ Abstract
Evaluating autonomous vehicles with controllability allows for scalable testing in counterfactual or structured settings, improving both efficiency and safety. We introduce LangTraj, a language-conditioned scene-diffusion model that simulates the joint behavior of all agents in traffic scenarios. By conditioning on natural language inputs, LangTraj enables flexible and intuitive control over interactive behaviors, generating nuanced and realistic scenarios. Unlike prior approaches that rely on domain-specific guidance functions, LangTraj incorporates language conditioning during training for more intuitive traffic simulation control. In addition, we propose a novel closed-loop training strategy for diffusion models to enhance realism in closed-loop simulation. To support language-conditioned simulation, we develop a scalable pipeline for annotating agent-agent interactions and single-agent behaviors, which we use to develop InterDrive, a large-scale dataset offering diverse and interactive labels for training language-conditioned diffusion models. Validated on the Waymo Motion Dataset, LangTraj demonstrates strong performance in both realism, language controllability, and language-conditioned safety-critical simulation, establishing a new paradigm for flexible and scalable autonomous vehicle testing.
Figure 1: Unconditioned (top), Text-Conditioned (mid), and Safety-Critical + Text-Conditioned (bot) Simulations by LangTraj. Adverarial collision guidance is applied in conjunction with text conditioning to generate the safety-critical scenarios shown in the bottom row. Language annotations are from the test set of InterDrive.
@@ 1 Introduction
Traditionally, structured testing of AVs has relied on manually designed scenarios to simulate failure cases or counterfactual situations, such as collisions or typical interactive behaviors. While effective for targeted testing, this approach is inherently limited in scalability. Recent advances in diffusion-based generative models demonstrate strong capabilities in simulating complex distributions with flexible controllability [39, 2], though this often depends on domain-specific heuristic guidance functions constructed based on human knowledge post-training. Our key insight is to leverage language conditioning during training to directly learn from the data distribution, enabling users—including non-experts—to generate counterfactual scenarios with ease. By conditioning on natural language inputs, we can significantly expand the range of possible scenarios, creating a more capable model as the data scales and allowing for diverse driving behaviors. Note that direct conditioning on language is orthogonal to guidance functions, allowing us to combine the strengths of both human knowledge and data-driven insights for more robust simulations.
@@
Furthermore, to support this modeling approach, we introduce InterDrive, a comprehensive dataset with 150k human-labeled annotations, specifically focusing on interactive agent-agent behaviors (e.g., merging, yielding, and passing). In addition, heuristic-based labels are used for single-agent behaviors such as lane changes and stops. This dataset ensures high-quality annotations for interaction modeling, providing the necessary diversity and richness for training models that capture both interactive and individualistic driving behaviors.
@@
  • InterDrive Dataset. We present InterDrive, a unique dataset of human-labeled interactive traffic behaviors, complemented by a secondary dataset of heuristically labeled single-agent behaviors. Together, these datasets support the training of models that capture a wide spectrum of traffic behaviors.
@@ 3 InterDrive Dataset
Natural language conditioning for traffic simulation relies heavily on the scale and quality of the underlying data. To address this, we introduce InterDrive, a central contribution of this work, designed to capture nuanced agent-agent interactions in real-world driving contexts. Our dataset is constructed through an extensive human-labeling effort on the Waymo Motion [3, 4] and NuPlan [1, 14] datasets, with additional single-agent behavioral labels generated via heuristic annotations.
@@ 3.1 Dataset Composition
The InterDrive dataset includes annotations from both the Waymo [3, 4] and NuPlan [1, 14] datasets, covering a diverse range of environments, agent types, and interaction scenarios. For Waymo, 125k scenes were annotated with human-generated interaction labels and 405k scenes annotated with heuristic labels. For nuPlan, 12k scenes were annotated with interaction labels and 150k scenes annotated with heuristic labels, each lasting 15 seconds.
@@ 3.2 Human-Labeling Process
    Classify Interaction Subtypes: For each chosen interaction type, labelers provide a more granular interaction subtype, capturing specifics like “Changing lane for overtaking,” and “Intersection yielding”. A full list of interaction subtypes is provided in Appendix D and in Fig. 2.
@@ 4 Problem Formulation
Each agent shares context $\mathbf{c}_{t}$, including a static HD map $I$ and historical states $\mathbf{s}_{t-T_{\text{hist}}:t}$. $g_{\theta}$ generates future trajectories $\{\mathbf{s}^{i}_{t:t+T}\}_{i=1}^{N}$ based on $(\mathbf{c}_{t},\mathbf{e}_{\text{lang}})$, trained on real-world data to ensure realism and adaptability to user-defined scenarios.
@@
$p_{\theta}(\tau_{k-1}|\tau_{k},\mathbf{c},\mathbf{e}_{\text{lang}}):=\mathcal{N}(\tau_{k-1};\mu_{\theta}(\tau_{k},k,\mathbf{c},\mathbf{e}_{\text{text}}),\Sigma_{k}).$
@@ 5 LangTraj
We introduce LangTraj, a scene-diffusion model designed to capture the joint distribution of interactive behaviors among all agents within multi-agent environments. To achieve this, LangTraj comprises two key components: 1) an encoder that effectively represents the scene context, including map features and historical agent behaviors, and 2) a denoiser capable of jointly predicting future agent behaviors while providing flexibility and user-driven control via natural language inputs. LangTraj supports trajectory generation conditioned on textual prompts and diffusion guidance, ensuring both flexibility and precise control. Additionally, we propose a novel algorithm for closed-loop training of diffusion models, further enhancing the model’s performance and applicability.
@@ 5.1.1 Scene Encoder
Inspired by [40, 26], we adopt a query-centric approach combined with graph neural networks (GNNs) to model spatiotemporal relationships in the scene. A key component of this approach is the attention mechanism, which encodes relative spatial and temporal information, capturing interactions and dependencies between scene elements. The encoder operates in a scene-independent local coordinate system. Each scene element extracts features within its own local reference frame, independent of global coordinates or the ego vehicle’s position. This formulation allows for symmetric encoding across agents scenarios without being affected by variations in absolute positioning. Through this process, the encoder produces an embedding $\mathbf{z}^{i}_{\text{enc}}=E_{enc}(I,\mathbf{S}_{t-T_{\text{hist}}:t})$ for each agent $i$.
@@ 5.1.2 Language Encoder
The Language Encoder module extracts agent-specific embeddings from natural language inputs and integrates them with the scene’s spatiotemporal context (see Fig. 3). This module serves two key functions: 1) providing explicit agent-specific conditioning and 2) encoding spatiotemporal data to define agent interactions.
@@
Our design is flexible regarding the choice of language model; following the practice of [25], we use DistillBERT [24], a smaller language encoder, to extract the $<$cls$>$ token embedding as a summary of the sentence, reducing computation cost. Empirically, we find this sufficient for language conditioning. In practice, we train the language encoder end-to-end with the diffusion model using LoRA (Low-Rank Adaptation), enabling parameter-efficient fine-tuning.
@@ 5.1.3 Denoiser
9:         Generate $M$ denoised candidates $\hat{s}^{(1:M)}$, select $\hat{s}^{(m^{*})}$ via $\arg\min_{m}D(\hat{s}^{(m)},s_{\text{target}})$
10:         Execute $\hat{s}^{(m^{*})}$, update state, compute loss $\mathcal{L}=\|\hat{s}^{(m^{*})}_{0:T}-s_{0:T}\|^{2}$, update $\theta$
11:     end for
@@
13:return $\theta$
@@ 5.3 Controllable Behavior via Diffusion Guidance
The details of the adopted guidance functions can be found in Appendix C. For classifier-free guidance, we interpolate between conditional and unconditional predictions from $g$:
@@ 6 Experimental Results
Table 1: Evaluation of InterDrive To enable a fair comparison with ProSim, we train LangTraj on the ProSim-Instruct-520k training set and evaluate its performance against the publicly released ProSim model using the ProSim-Instruct-520k test set.
@@ 6.1 Preliminaries
We validate the efficacy of our proposed framework through experiments on real-world driving data from the Waymo Open Motion Dataset (WOMD) [5]. Our results demonstrate that the framework can generate realistic and controllable behaviors that align closely with interactive language descriptions. We train LangTraj on the training splits of InterDrive and ProSim-Instruct-520k, which we describe in detail in Appendix A.
@@
Table 2: Evaluation of LangTraj. Closed-loop evaluation on ProSim-Instruct-520k test set with and without text conditioning. Both methods trained on ProSim-Instruct-520k train set.
@@
Our goal is to validate that LangTraj generates interactive behaviors that are both realistic and controllable using InterDrive. We evaluate model performance across two main dimensions: realism (denoted ”Meta” in the tables) and controllability.
For realism, we follow the Waymo Open Sim Agent Challenge (WOSAC), which evaluates the distributional realism of generated trajectories [18]. The challenge assesses how well joint simulation rollouts recover held-out ground truth behavior across multiple aspects, including kinematics, agent interactions, and map adherence. Further details can be found in Sec. A.4.
To evaluate controllability, we condition the model on ground-truth future behavior descriptions and measure the improvement in minADE (m) compared to the unconditional model. This comparison quantifies the model’s ability to align generated behaviors with user-specified input when provided with accurate behavior descriptions.
@@ 6.2 Evaluation of InterDrive
We train separate instantiations of LangTraj on the training sets of InterDrive and ProSim-Instruct-520k, performing closed-loop evaluations both with and without text conditioning on their respective test sets. Results summarized in Tab. 2 highlight the effectiveness of LangTraj in enhancing simulation realism via language conditioning. Specifically, on InterDrive, incorporating text descriptions maintains the meta-realism metric at 0.72, accompanied by a notable 13.6% reduction in minADE. This indicates that our model consistently generates trajectories that are more realistic and closely aligned with the ground truth. A similar improvement is observed when evaluating on ProSim-Instruct-520k [31], where text conditioning results in a 12.80% reduction in minADE, confirming that language inputs consistently enhance simulation quality across diverse datasets.
@@ 6.4 Text-Conditioned Safety-Critical Simulation
We demonstrate LangTraj’s ability to extend beyond the training distribution by generating safety-critical scenarios conditioned on text inputs using guided sampling, as illustrated in Tab. 4 and Fig. 1. Given the focus on collision-prone situations, standard interactive metrics are less informative; thus, we instead evaluate simulation quality using kinematics and map adherence metrics. Our findings indicate that LangTraj successfully generates realistic traffic scenarios as well as targeted, rare, safety-critical events, achieving collision rates as high as 40%. Additionally, integrating text conditioning enhances map adherence even under collision-focused guidance, highlighting a promising approach that combines explicit guidance with textual inputs to improve scenario diversity and simulation control. Supplementary videos are available for readers to explore further demonstrations of combining text-based conditioning and collision guidance to produce diverse critical scenarios.
6.5 Waymo Sim Agents Challenge
We evaluate LangTraj on the Waymo Sim Agents Challenge (WOSAC) benchmark test set (unconditioned simulation), with results presented in Tab. 3. Compared to state-of-the-art diffusion models, LangTraj performs competitively among the best diffusion-based approaches, such as VBD [9] and SceneDiffuser [13]. LangTraj achieves higher Map metrics than both VBD and SceneDiffuser, demonstrating improved adherence to road constraints based on closed-loop training, as discussed in Sec. 6.6. However, kinematic metrics remain an area for future improvement, suggesting potential refinements in generating higher likelihood interms of speed and acceleration. Note that unlike SceneDiffuser, which applies generalized hard constraints such as on-road or collision avoidance during sampling or other methods that incorporate guidance, LangTraj generates trajectories purely from its learned distribution.
Guidance Text Collision Rate $\uparrow$ Kinematic $\uparrow$ Map $\uparrow$ mADE $\downarrow$
@@ 6.5 Waymo Sim Agents Challenge
Collision $\times$ 0.41 0.39 0.70 4.93
Collision $\checkmark$ 0.43 0.41 0.74 4.43
@@ 7 Conclusion
A key component of our approach is InterDrive, which focuses on interactive agents, providing rich human-labeled and heuristic annotations that enhance realism and diversity. By prioritizing agent-agent interactions, InterDrive strengthens training for language-conditioned models. Empirical results on the Waymo Motion Dataset confirm that LangTraj generates realistic trajectories aligned with language instructions while maintain simulation realism. In summary, LangTraj demonstrates the potential of language-conditioned diffusion models for AV simulation, integrating natural language inputs with guidance-driven scenario design.

### v2

@@ Abstract
Evaluating autonomous vehicles with controllability allows for scalable testing in counterfactual or structured settings, improving both efficiency and safety. We introduce LangTraj, a language-conditioned scene-diffusion model that simulates the joint behavior of all agents in traffic scenarios. By conditioning on natural language inputs, LangTraj enables flexible and intuitive control over interactive behaviors, generating nuanced and realistic scenarios. Unlike prior approaches that rely on domain-specific guidance functions, LangTraj incorporates language conditioning during training for more intuitive traffic simulation control. In addition, we propose a novel closed-loop training strategy for diffusion models to enhance realism in closed-loop simulation. To support language-conditioned simulation, we develop a scalable pipeline for annotating agent-agent interactions and single-agent behaviors, which we use to develop InterDrive, a large-scale dataset offering diverse and interactive labels for training language-conditioned diffusion models. Validated on the Waymo Open Motion Dataset, LangTraj demonstrates strong performance in both realism, language controllability, and language-conditioned safety-critical simulation, establishing a new paradigm for flexible and scalable autonomous vehicle testing. Project website: https://langtraj.github.io/.
Figure 1: Unconditioned (top), Text-Conditioned (mid), and Safety-Critical + Text-Conditioned (bot) Simulations by LangTraj. Adversarial collision guidance is applied in conjunction with text conditioning to generate the safety-critical scenarios shown in the bottom row, where the dark blue car serves as the adversarial agent. Language annotations are from the test set of InterDrive.
@@ 1 Introduction
Traditionally, structured testing of AVs has relied on manually designed scenarios to simulate failure cases or counterfactual situations, such as collisions or typical interactive behaviors. While effective for targeted testing, this approach is inherently limited in scalability. Recent advances in diffusion-based generative models demonstrate strong capabilities in simulating complex distributions with flexible controllability [35, 2], though this often depends on domain-specific heuristic guidance functions constructed based on human knowledge post-training. Our key insight is to leverage language conditioning during training to directly learn semantics from the data distribution, enabling users—including non-experts—to generate counterfactual scenarios with ease. By conditioning on natural language, we can significantly expand the range of possible scenarios, creating a more capable model as the data scales and allowing for diverse driving behaviors. Note that direct conditioning on language is orthogonal to guidance functions, allowing us to combine the strengths of both human knowledge and data-driven insights for more diverse simulations.
@@
To enable language-conditioned simulation, we introduce InterDrive, a comprehensive dataset with 150k human-labeled annotations, specifically focusing on interactive agent-agent behaviors (e.g., merging, yielding, and passing) with rich language annotations. In addition, we include heuristic-based labels for single-agent behaviors such as directional intent and lane changes. This dataset ensures high-quality annotations for interaction modeling, providing the necessary diversity and richness for training models that capture both interactive and individualistic driving behaviors.
@@
  • InterDrive Dataset. We present InterDrive, a new dataset of 150k human-labeled interactive traffic scenarios, supplemented with heuristically labeled single-agent behaviors, enabling scalable training of language-conditioned simulation models.
@@ 3 InterDrive Dataset
Natural language conditioning for traffic simulation relies heavily on the scale and quality of the underlying data. To address this, we introduce InterDrive, a central contribution of this work, designed to capture nuanced agent-agent interactions in real-world driving contexts. Our dataset is constructed through an extensive human-labeling effort on the Waymo Motion [3] and NuPlan [1] datasets, with additional single-agent behavioral labels generated via heuristic annotations.
@@ 3.1 Dataset Composition
The InterDrive dataset includes annotations from both the Waymo [3] and NuPlan [1] datasets, covering a diverse range of environments, agent types, and interaction scenarios. For Waymo, 125k scenes were annotated with human-generated interaction labels and 405k scenes annotated with heuristic labels. For nuPlan, 12k scenes were annotated with interaction labels and 150k scenes annotated with heuristic labels, each lasting 15 seconds.
@@ 3.2 Human-Labeling Process
    Classify Interaction Subtypes: For each chosen interaction type, labelers provide a more granular interaction subtype, capturing specifics like “Changing lane for overtaking,” and “Intersection yielding”. A full list of interaction subtypes is provided in Appendix E and in Fig. 2.
@@ 4 Problem Formulation
Each agent shares context $\mathbf{c}_{t}$, including HD map $I$ and historical states $\mathbf{S}_{t-T_{\text{hist}}:t}$. The function $g_{\theta}$ generates future trajectories $\{\mathbf{s}^{i}_{t:t+T}\}_{i=1}^{N}$ based on $(\mathbf{c}_{t},\mathbf{e}_{\text{lang}})$, trained on real-world data to ensure realism and flexibility to user-defined scenarios.
@@
$p_{\theta}(\tau_{k-1}|\tau_{k},\mathbf{c},\mathbf{e}_{\text{lang}}):=\mathcal{N}(\tau_{k-1};\mu_{\theta}(\tau_{k},k,\mathbf{c},\mathbf{e}_{\text{lang}}),\Sigma_{k}).$
@@ 5 LangTraj
We introduce LangTraj, a scene-diffusion model designed to capture the joint distribution of interactive behaviors among all agents within multi-agent environments. To achieve this, LangTraj comprises two key components: 1) an encoder that effectively represents the scene context, including map features and historical agent behaviors, and 2) a denoiser capable of jointly predicting future agent behaviors while providing flexibility and user-driven control via natural language inputs. LangTraj supports trajectory generation conditioned on textual prompts and diffusion guidance, ensuring both flexibility and precise control. Additionally, we propose a novel algorithm for closed-loop training of diffusion models, further enhancing the model’s performance and applicability in closed-loop simulation settings.
@@ 5.1.1 Scene Encoder
Inspired by [36, 22], we adopt a query-centric approach combined with Graph Neural Networks (GNNs) to model spatiotemporal relationships in the scene. A key component of this approach is the attention mechanism, which encodes relative spatial and temporal information, capturing interactions and dependencies between scene elements. The encoder operates in a scene-independent local coordinate system. Each scene element extracts features within its own local reference frame, independent of global coordinates or the ego vehicle’s position. This formulation allows for symmetric encoding across agents without being affected by variations in absolute positioning. Through this process, the encoder produces an embedding $\mathbf{z}^{i}_{\text{enc}}=E_{enc}(I,\mathbf{S}_{t-T_{\text{hist}}:t})$ for each agent $i$.
@@ 5.1.2 Language Encoder
The Language Encoder module extracts agent-specific embeddings from natural language inputs and integrates them with the scene’s spatiotemporal context (see Fig. 3). This module serves two key functions: 1) providing explicit agent-specific conditioning and 2) encoding spatiotemporal data to capture agent interactions.
@@
Our design is flexible regarding the choice of language model; following the practice of [21], we use DistillBERT [20], a smaller language encoder, to extract the <<cls>> token embedding as a summary of the sentence, reducing computation cost. Empirically, we find this sufficient for language conditioning. In practice, we train the language encoder end-to-end with the diffusion model using LoRA (Low-Rank Adaptation), enabling parameter-efficient fine-tuning.
@@ 5.1.3 Denoiser
9:   Generate $M$ denoised candidates $\hat{s}^{(1:M)}$
10:   Select $\hat{s}^{(m^{*})}$ via $\arg\min_{m}D(\hat{s}^{(m)},s_{\text{target}})$
11:   Execute $\hat{s}^{(m^{*})}$, update state
@@
13:  Compute loss $\mathcal{L}=\|\hat{s}_{0:T}-s_{0:T}\|^{2}$, update $\theta$
14:end for
15:return $\theta$
@@ 5.3 Controllable Behavior via Diffusion Guidance
The details of the adopted guidance functions can be found in Appendix D. For classifier-free guidance, we interpolate between conditional and unconditional predictions from $g$:
@@ 6 Experimental Results
Table 1: Evaluation of InterDrive: We train LangTraj separately on the InterDrive and ProSim-Instruct-520k training sets and evaluate its performance on their respective test sets.
@@ 6.1 Preliminaries
We validate our framework through experiments on real-world driving data from the Waymo Open Motion Dataset (WOMD) [3]. We train LangTraj on the training splits of InterDrive and ProSim-Instruct-520k, described in Appendix A, and evaluate its ability to generate behaviors that are both realistic and controllable.
@@
Table 2: Evaluation of LangTraj. Closed-loop evaluation on the ProSim-Instruct-520k test set, with and without text conditioning. We compare the performance of our method (LangTraj) against the publicly released ProSim model, both trained on the ProSim-Instruct-520k training set, ensuring a fair evaluation.
@@
For realism (denoted ”Meta” in the tables), we follow the Waymo Open Sim Agent Challenge (WOSAC), which evaluates the distributional realism of generated trajectories [14]. The challenge assesses how well joint simulation rollouts recover held-out ground truth behavior across multiple aspects, including kinematics, agent interactions, and map adherence. Further details can be found in Sec. A.5. To evaluate controllability, we condition the model on ground-truth future behavior descriptions and measure the improvement in minADE compared to the unconditional model. This comparison quantifies the model’s ability to align generated behaviors with user-specified input when provided with behavior descriptions.
@@ 6.2 Evaluation of InterDrive
We train separate instantiations of LangTraj on the training sets of InterDrive (InterDrive) and ProSim-Instruct-520k, performing closed-loop evaluations with and without text conditioning. As shown in Tab. 2, language conditioning on InterDrive yields a 13.6% reduction in minADE while maintaining a meta-realism score of 0.72, indicating that our language annotations provide useful supervisory signals. A similar improvement is observed on ProSim-Instruct-520k (12.8% minADE reduction). Notably, InterDrive focuses on interactive agent-agent behaviors, whereas ProSim-Instruct primarily consists of single-agent maneuvers.
Text Conditioning Meta $\uparrow$ Kinematic $\uparrow$ Interactive $\uparrow$ Map $\uparrow$ mADE $\downarrow$
None 0.72 0.41 0.80 0.80 2.65
Direct Condition (Ours) 0.72 0.42 0.80 0.79 2.29
LLM-Based Guidance (CTG++ Style) 0.70 0.42 0.75 0.80 2.70
Table 4: Text Conditioning Evaluation. We compare our proposed direct text conditioning method to the LLM-based guidance method proposed by CTG++  [34].
6.3 LLM-based Guidance v.s Direct Conditioning
In Tab. 4, we compare our proposed direct language conditioning method with the LLM-based guidance approach from CTG++ [34], which uses GPT to generate differentiable code-based guidance functions from language prompts (see Sec. A.3 for details). On InterDrive scenarios, we find that LLM-based guidance leads to worse alignment with the text prompt—even underperforming the no-text baseline. In contrast, direct conditioning produces behaviors more aligned with the ground truth (13.6% reduction) while preserving simulation realism. Direct conditioning also outperforms LLM-based guidance in counterfactual settings, as discussed in Sec. 6.5. While LLM-based guidance is effective for enforcing constraints like collision avoidance and on-road driving [34], it still struggles with interaction-level behavior descriptions, underscoring the benefit of direct language grounding.
@@ 6.4 Evaluation of LangTraj
Beyond instruction-following tasks, we evaluate LangTraj on the Waymo Sim Agents Challenge Benchmark (unconditioned simulation), with results presented in Tab. 3. Compared to state-of-the-art diffusion models, LangTraj performs competitively among the best diffusion-based approaches, such as VBD [7] and SceneDiffuser [10].
@@ 6.5 Text-Conditioned Safety-Critical Simulation
We demonstrate LangTraj’s ability to extend beyond the training distribution by generating safety-critical scenarios conditioned on text inputs using guided sampling, as illustrated in Tab. 5 and Fig. 1. Given the focus on collision-prone situations, standard interactive metrics are less informative; thus, we instead evaluate simulation quality using kinematics and map adherence metrics. Our findings indicate that LangTraj successfully generates realistic traffic scenarios as well as targeted, rare, safety-critical events, achieving collision rates as high as 40%. Compared to LLM-based guidance, LangTraj achieves a 10% higher collision rate and outperforms CTG++ across all metrics in safety-critical situations. Additionally, integrating direct text conditioning enhances map adherence even under collision-focused guidance, highlighting a promising approach that combines explicit guidance with textual inputs to improve scenario diversity and simulation control. Qualitative examples are provided in the supplementary videos.
COLLISION
Text Conditioning Guidance Rate $\uparrow$ Kinematic $\uparrow$ Map $\uparrow$ mADE $\downarrow$
@@
None $\checkmark$ 0.41 0.39 0.70 4.93
Direct Condition (Ours) $\checkmark$ 0.43 0.41 0.74 4.43
LLM-based Guidance (CTG++ Style) $\checkmark$ 0.33 0.37 0.72 4.67
@@ 7 Conclusion
A key component of our approach is InterDrive, which focuses on interactive agents, providing rich human-labeled and heuristic annotations that enhance realism and diversity. By prioritizing agent-agent interactions, InterDrive strengthens training for language-conditioned models. Empirical results on the Waymo Motion Dataset show that LangTraj generates realistic, language-aligned trajectories while preserving simulation realism. In summary, LangTraj demonstrates the potential of language-conditioned diffusion models for AV simulation by combining natural language with guidance-based control.
Acknowledgements
This work was part of W.J. Chang’s summer internship at NEC Labs America, and he is also supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 2146752. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. This study was funded in part by the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government via the Hong Kong Centre for Logistics Robotics. The authors would like to thank Yichen Xie and Rian Tian for their insightful discussions, and Chih-Ling Chang for her helpful suggestions and assistance with figures and presentations.